# A note on the triple product property for finite groups with abelian normal subgroups of prime index

### v2

@@ 3. Technical Results
Let $G$ and $p$ be as given. Let $H\unlhd G$ be abelian of index $p$. Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $H_{0}=H,H_{1},\ldots,H_{p-1}$ of $H$ in $G$. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$. By Observation 3.1 part (2) $\overline{S},\overline{T},\overline{U}\leq G/H$, and if $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4), and $S_{0}T_{0}U_{0}\mathrel{\unlhd}H$ (as $S_{0}T_{0}U_{0}\leq H$ and $H$ is abelian) and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2).
@@
Assume $|S||T||U|=\sigma\tau\upsilon|S_{0}||T_{0}||U_{0}|>|G|=p|H|$. Clearly at least two of $\sigma,\tau,\upsilon$ exceed $1$, that is, are equal to $p$. Without loss of generality we may assume $\sigma=\tau=p$. If $\upsilon=1$ then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{|H|}{p}$, and this implies that $|S_{0}||T_{0}||U_{0}|=|H|$ , that is, $S_{0}T_{0}U_{0}=H$. However, $S$ and $T$ are supported on all $p$ cosets of $H$ in $G$, and using Lemma 3.3 any non-trivial coset $xH$ can be chosen such that $S_{x}:=S\cap xH\neq\emptyset$, $T_{x}:=T\cap xH\neq\emptyset$ and $S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$, which is a contradiction. Thus $\sigma=\tau=\upsilon=p$ and $S,T,U\not\leq H$.
@@
So the assumption $|S||T||U|>|G|=p|H|$ implies that $S_{0}T_{0}U_{0}$ is a subgroup of $H$ of order $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}\leq|H|$. Furthermore applying part (1) of Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are collections of cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. As $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$. Now applying part (2) of Lemma 3.3 it follows that $\{S_{x}^{-1}T_{x}U_{0}\}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, and similarly that $T_{0}\trianglelefteq G$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, these collections only intersecting trivially in $S_{0}T_{0}U_{0}$. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.

### v5

@@ 4. The Main Result
Let $G$ and $p$ be as given. Let $H\unlhd G$ be abelian of index $p$. Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $x_{0}H=H,x_{1}H,\ldots,x_{p-1}H$ of $H$ in $G$ with representatives $x_{0}=1,x_{1},\ldots,x_{p-1}\in G$ respectively. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$ as defined in Observation 3.1. By part (2) of that Observation $\overline{S},\overline{T},\overline{U}\leq G/H$, and if $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. The sets $S_{x},T_{y},U_{z}$, as defined in Lemma 3.3 will also be used at certain points: these are defined as $S_{x}:=S\cap xH$, $T_{y}:=T\cap yH$, $U_{z}:=U\cap zH$ for any $xH,yH,zH\in G/H$. By definition, if $xH\in G/H$ then $xH\in\overline{S}\iff S_{x}\neq\emptyset$, and analagously for $T_{y}$ and $U_{z}$.
Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4), and $S_{0}T_{0}U_{0}\mathrel{\unlhd}H$ (as $S_{0}T_{0}U_{0}\leq H$ and $H$ is abelian) and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2).
@@
Suppose $|S||T||U|=\sigma\tau\upsilon|S_{0}||T_{0}||U_{0}|>|G|=p|H|$. Then at least two of $\sigma,\tau,\upsilon$ must exceed $1$, that is, must be equal to $p$. Without loss of generality let $\sigma=\tau=p$, which means that $S$ and $T$ are supported on all $p$ cosets of $H$, or, equivalenty, $\overline{S}=\overline{T}=G/H$ and all sets $S_{x}$ and $T_{y}$ as defined above are non-empty. Suppose $\upsilon=1$, that is, $U=U_{0}$. Then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{|H|}{p}$, i.e. $\frac{|H|}{|S_{0}||T_{0}||U_{0}|}=|H:S_{0}T_{0}U_{0}|<p$, which implies that $S_{0}T_{0}U_{0}=H$ and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=|H|$, and, using Lemma 3.3, elements $s_{x}\in S_{x}\neq S_{0}$, $t_{x}\in T_{x}\neq T_{0}$ can be chosen so that the translate $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$, a contradiction.
@@
So, if $|S||T||U|>|G|=p|H|$ then $\sigma=\tau=\upsilon=p$ and $S,T,U\not\leq H$ and $|S_{0}T_{0}U_{0}|<|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}<|H|$. Also, as $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$ (and in the same way $T_{0}=T\cap H\trianglelefteq\langle T,H\rangle=G$). Now applying Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are each collections of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.
A simple corollary is noted for some special cases,
Corollary 4.2.
If $G$ is a group with an abelian normal subgroup of prime index $p$ such that $(2p-1)\nmid|G|$, that is, (2p1)(2p-1) does not divide $|G|$, then $\rho_{0}(G)\leq\frac{1}{2}p$. If $G$ is a $p$-group with an abelian subgroup of index $p$ then $\rho_{0}(G)=1$.
Proof.
In the first case, if $H\leq G$ is abelian and normal of prime index $p$ such that $(2p-1)\nmid|G|$ then $(2p-1)\nmid|H|$, and if $(S,T,U)$ is a subgroup TPP triple of $G$ then the subgroup $S_{0}T_{0}U_{0}<H$, as defined above, cannot be as large as $\frac{|H|}{(2p-1)}$, i.e. $|S_{0}||T_{0}||U_{0}|<\frac{|H|}{(2p-1)}$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{2p}=\frac{1}{2}p|G|$. In the second case, if $G$ is a $p$-group then the product of the orders of any of its subgroups is a $p$-power, and if $G$ has an abelian subgroup of index $p$ (which is thus necessarily normal in $G$), and $(S,T,U)$ is a subgroup TPP triple of $G$, then $|S||T||U|$ must be a $p$-power such that $|S||T||U|<\frac{p^{2}}{2p-1}|G|<p|G|$, i.e. $|S||T||U|\leq|G|$. ∎
The contrapositive of this corollary explains the non-trivial $\rho_{0}$ values computed by Hedtke and Murthy for all the groups listed in [5, Tables 1-4] which satisfy the conditions of the corollary while achieving $\rho_{0}>1$ or $\rho_{0}>\frac{1}{2}p$, and this has been verified by GAP computations in these groups.
5. Analysis and Conclusion
In relation to the more general notion of TPP ratio $\rho$ the theorem means that in groups with abelian normal subgroups of prime index $p$ no three subgroups can realise the TPP with a triple size exceeding $\frac{p^{2}}{2p-1}|G|$. To achieve the latter one must look for triples of subsets at least one of which is not a subgroup.
In the case of $p=2$ and the dihedral groups $D_{2n}$ the theorem proves that $\rho_{0}(D_{2n})\leq\frac{4}{3}$ [5, Conjecture 7.5], and together with the data for TPP triples in groups of order $\leq 32$ obtained via exhaustive computational search [5, Table 1], strengthens the conjecture that $\rho\leq\frac{4}{3}$ for groups with cyclic (and normal) subgroups of index $2$ [5, Conjecture 7.6].
For $p>2$ the smallest known group containing a cyclic normal subgroup of index $p$ and a known $\rho$ is one of type $C_{7}\rtimes C_{3}$ (GAP ID [21, 1]), a non-abelian group of order $21=3\cdot 7$ containing a cyclic normal subgroup $C_{7}$ of index $p=3$, and which realises a largest TPP triple of type $(3,3,3)$ and size $3^{3}=27$, and satisfies the bound $\frac{27}{21}<\frac{3^{2}}{2\cdot 3-1}=\frac{9}{5}=1.8$ [5, Table 1]. The same data also shows that there are actually no groups of order $\leq 32$ containing cyclic normal subgroups of index $3$ and realising $\rho>\frac{9}{5}$. This seems to be true even for larger groups of order up to $100$, as indicated in the results of Xiang et. al. in 2018, who use a computational approach based on evolutionary search algorithms to look at TPP capacity $\beta$ (and other parameters related to group-theoretic matrix multiplication) in a selection of groups of order up to $100$ [8, Tables 1, 3-4]: some simple GAP computations show that of all groups listed in those tables which realise $\rho>\frac{9}{5}$ none contain cyclic normal subgroups of index $3$. Noting that a subgroup of a group which is of smallest prime index is always normal, this suggests the following conjecture.
Conjecture 5.1.
If $G$ is a group with a cyclic normal subgroup of prime index $p$, and $(S,T,U)$ is any TPP triple of $G$, then $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|$, i.e. $\rho(G)\leq\frac{p^{2}}{2p-1}$. For this bound to be as tight as possible $p$ can be taken to be the smallest such prime, in which case the cyclic subgroup of index $p$ will necessarily be normal.
The conjecture is false for the more general case of abelian normal subgroups of prime index: for $p=2$, where the upper bound to be respected is $\rho\leq\frac{4}{3}$, the smallest known counterexamples are two $2$-groups of order $32$, one of type $(C_{4}\times C_{4})\rtimes C_{2}$ (GAP ID [32, 11]) and another of type $(C_{2}\times C_{2}\times C_{2}\times C_{2})\rtimes C_{2}$ (GAP ID [32, 27]), both containing a non-cyclic abelian subgroup of index $2$, but no cyclic abelian subgroups of index $2$, and realising $\rho=1.5>\frac{4}{3}$ (via TPP triples of type $(6,4,2)$).